OpenStudy (adamaero):

Completing the square? or some other way? 2 + x^2 - x^(-2)

OpenStudy (anonymous):

\[2+x^2-\frac{1}{x^2}\] is what you have

OpenStudy (anonymous):

is it set equal to zero or sommat? this is not an equation, nothing to solve

OpenStudy (adamaero):

\[2+e ^{2t}-e ^{-2t}\]

OpenStudy (anonymous):

still not an equation, still nothing to solve

OpenStudy (anonymous):

really it might be best to post the whole thing

OpenStudy (adamaero):

\[\sqrt{2+e ^{2t}-e ^{-2t}}\]

OpenStudy (adamaero):

somehow that equates to \[\sqrt{(e ^{t} + e ^{-t})^2}\]

OpenStudy (anonymous):

sure does, multiply out and you will see it right away

OpenStudy (adamaero):

how was I suppose to derive that!?

OpenStudy (anonymous):

that i cannot answer

OpenStudy (adamaero):

waaaaa? but you're the best? and all you have for this one is trial and error??

OpenStudy (anonymous):

once you see it it is totally clear

OpenStudy (adamaero):

meh

OpenStudy (adamaero):

thanks tho

OpenStudy (anonymous):

\[(\frac{1}{x}+x)^2=\frac{1}{x^2}+2\times x\times \frac{1}{x}+x^2\]